OpenStudy (sgadi):

\[\int\limits_{0}^{\infty}\left(\frac{e^{-at}}{\sqrt{c+d\cos{e+t}}}\right)dt\]

OpenStudy (sgadi):

\[\int\limits_{0}^{\infty}\left(\frac{e^{-at}}{\sqrt{c+d\cos{e+t}}}\right)dt\]